Other Math

Given the following data of a quadrilateral: ∠ABC=104°, ∠BCD=89°, ∠ABD=60°, ∠ACD=53°, and AD=1000 m

1.)Determine the length of line BC (in meters, 2 decimal place):   

2.)Determine ∠CAD:

3.)Determine ∠BDA:



1
Expert's answer
2021-05-07T09:59:34-0400


Values to find: "x, y, BC" (see picture).

1) Obviously "x+y=80\degree". Via Sinus Theorem

"AB=\frac{AC \sin(36\degree)}{\sin(104\degree)}" & "AD=\frac{AC\sin(53\degree)}{\sin(37\degree+y)}" & "AB=\frac{AD\sin(y)}{\sin(60\degree)}" =>

"\frac{AC\sin(53\degree)\sin(y)}{\sin(37\degree+y)\sin(60\degree)}=\frac{AC\sin(36\degree)}{\sin(104\degree)}"

Using "\sin(53\degree)=\sin(90\degree-37\degree)=\cos(37\degree)" we can rewrite previous equation as

"\frac{\cos(37\degree)\sin(y)}{\sin(37\degree)\cos(y)+\cos(37\degree)\sin(y)}=\frac{\sin(60\degree)\sin(36\degree)}{\sin(104\degree)}" =>

"\frac{\sin(104\degree)}{\sin(60\degree)\sin(36\degree)}=1+\tg(37\degree)\ctg(y)" =>

"y=\arctg\biggr[\frac{1}{\tg(37\degree)}(\frac{\sin(104\degree)}{\sin(60\degree)\sin(36\degree)}-1)\biggl]"

"y \approx\arctg(1,19)\approx 40\degree"

"x=80\degree-y\approx 40\degree"

2) Via Sinus Theorem "AB=\frac{BC\sin(36\degree)}{\sin(40\degree)}" & "AB=\frac{AD\sin(y)}{\sin(60\degree)}" =>

"BC=\frac{AD\sin(y)\sin(40\degree)}{\sin(60\degree)\sin(36\degree)}"

"BC\approx 1000m*0,798=798m"